OpenStudy (anonymous):

Need Help Please? 5x-7/3-1=5

OpenStudy (anonymous):

5 x-10/3 = 5

OpenStudy (anonymous):

x = 5/3

OpenStudy (anonymous):

x = 11/5 x = 13/5 x = 4 x = 5

OpenStudy (anonymous):